# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v3

@@ 4.2 Experimental results
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.2, A4 and A5).
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.3, A6 and A7 in the Appendix. However, the trend w.r.t. $C$ is not as clear as with $S$.
(a) as function of $S$ (shots)
(b) as function of $C$ (classes) at $S=32$
[…]
RMIA In Figure 2 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr). More results for RMIA are in Figures A.6, A.7 and A.8 in the Appendix.
@@ 4.3 Model to predict dataset vulnerability
The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3. We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold & Perktold, 2010). The general form of the model can be found in Equation 27, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
@@
We utilise MIA results of ViT-B (Head) (see Table A4) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.4 and Table A9 in the Appendix. Figure 3 shows the parameters of the prediction model fitted to the training data. For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts. However, the coefficient value decreases for small fpr. This is perhaps because the power-law in Equation (24) only holds for large $S$, and for small fpr Equation (24) significantly underestimates the vulnerability in small-$S$ regime (see Appendix D).
Figure 3: Coefficient values for different fpr when fitting a regression model based on Equation 27 fitted on data from ViT-B (Head) with LiRA (Table A4). The error bars display the 95% confidence intervals based on Student’s t-distribution.
@@
  • R-50 (FiLM): Figure 4(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section E.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A8 in the Appendix.
@@
(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al. (2023), see Table A8) 2 right: trained from scratch (data from Carlini et al. (2022)) . The linear model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted. Note that the results from Carlini et al. (2022) use data augmentation while we do not.
@@ 5 Discussion
Under the GDPR and similar legal regimes, machine learning (ML) models trained on personal data that memorise the data are personal data and need to be carefully protected. Our work analyses in which cases trained models would most likely be personal data and in which cases they might not be. This will help in evaluating the risk of different kinds of models, favouring less risky models when possible and paying extra attention to more risky cases.
As the best means of protecting privacy, differential privacy, reduces the utility of models, it is important to understand when it is necessary. Our results highlight that models are most vulnerable to MIA when the number of examples per class is low. This reinforces the well-known unfortunate dilemma: differential privacy seems most important in the regime where it likely reduces the model utility the most.
One major reason for MIA vulnerability can be memorisation of the training data. Feldman & Zhang (2020) experimentally test memorisation in neural network training, and find that according to their definition, a large fraction of the training data are memorised when training from scratch, while only few are when fine-tuning. This is aligned with our results that indicate from scratch training to be more vulnerable than fine-tuning.
Our model singles out the number of examples per class $S$ as the key predictor of model vulnerability. While our results are based on balanced datasets, it seems likely that minority classes and outliers would be more vulnerable as predicted by our model given their smaller $S$. Studying the vulnerability in unbalanced data and in the presence of outliers is an important area of future work.
@@
Our work measures the vulnerability through running two particular MIAs. There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attacks. Formal bounds on attack susceptibility would require something like differential privacy. Both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable.
@@ Acknowledgments
Appendix A Formulating LiRA and RMIA for Section 3
Let $\mathcal{M}$ be our target model and $\ell(\mathcal{M}(x),y_{x})$ be the loss of the model on a target example $(x,y_{x})$. The goal of MIA is to determine whether $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$. This can be formulated as a hypothesis test:
$H_{0}:(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$ (A1)
$H_{1}:(x,y_{x})\in\mathcal{D}_{\mathrm{target}}.$ (A2)
A.1 LiRA
Following (Carlini et al., 2022), we formulate the Likelihood Ratio Attack (LiRA). LiRA exploits the difference of losses on the target model under $H_{0}$ and $H_{1}$. To model the IN/OUT loss distributions with few shadow models, LiRA employs a parametric modelling. Particularly, LiRA models $t_{x}$ by a normal distribution. That is, the hypothesis test formulated above can be rewritten as
$H_{0}:t_{x}\sim\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})$ (A3)
$H_{1}:t_{x}\sim\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2}).$ (A4)
@@ A.1 LiRA
$\mathrm{LR}(x)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x)^{2},\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}.$ (A5)
LiRA rejects $H_{0}$ if and only if
$\mathrm{LR}(x)\geq\beta,$ (A6)
concluding that $H_{1}$ is true, i.e., identifying the membership of $(x,y_{x})$. Thus, the true positive rate of this hypothesis test given as
$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (A7)
@@
A.2 RMIA
By modelling $t_{z}$ by a normal distribution, Zarifzadeh et al. (2024) approximate the pairwise likelihood ratio as
$\mathrm{LR}(x,z)=\frac{p(\theta\mid x,y_{x})}{p(\theta\mid z,y_{z})}\approx\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{x,z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{z,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z,z}^{2})},$ (A8)
where $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ are, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models when the training set contains $a$ but not $b$. Then RMIA exploits the probability of rejecting the pairwise likelihood ratio test over $(z,y_{z})\sim{\mathbb{D}}$:
$\mathrm{Score}_{\mathrm{RMIA}}(x)=\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right).$ (A9)
Thus, RMIA rejects $H_{0}$ if and only if
$\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta,$ (A10)
identifying the membership of $x$. Hence the true positive rate of RMIA is given as
$\textsc{tpr}_{\mathrm{RMIA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\right),$ (A11)
Appendix B On the assumptions in Section 3
B.1 The assumption of shared variances
@@
In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$, and that for RMIA $\sigma_{x}=\sigma_{x,x}=\sigma_{z,x}$, $\sigma_{z}=\sigma_{x.z}=\sigma_{z,z}$, $\hat{\sigma}_{x}=\hat{\sigma}_{x,x}=\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{z}=\hat{\sigma}_{x.z}=\hat{\sigma}_{z,z}$. Utilising the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
From the proof of Theorem 3 (see Section C.3), the LiRA scores are
$s_{y_{x}}^{\mathrm{(in)}}$ $=\langle x,\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{y_{x},i}+x\right)\rangle=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle+\langle x,\frac{1}{S}(x-x_{y_{x},S})\rangle$ (A12)
$s_{y_{x}}^{\mathrm{(out)}}$ $=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle,$ (A13)
@@
respectively, when $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$ and when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$. Then, since for the simplified model scores follow Gaussian distributions, we have
@@
$\sigma_{\mathrm{in}}^{2}$ $=\hat{\sigma}_{\mathrm{in}}^{2}=\operatorname{Var}(s_{y_{x}}^{\mathrm{(in)}})=\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}$ (A14)
$\sigma_{\mathrm{out}}^{2}$ $=\hat{\sigma}_{\mathrm{out}}^{2}=\operatorname{Var}(s_{y_{x}}^{\mathrm{(out)}})=\frac{1}{S}||x||^{2}s^{2}.$ (A15)
Thus, the differences $\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}-\hat{\sigma}_{\mathrm{out}}$ are negligible for large $S$. Similarly, we have
$\sigma_{x,x}^{2}$ $=\hat{\sigma}_{x,x}^{2}=\operatorname{Var}(s_{y_{x}}^{(x)}(x))=\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}$ (A16)
$\sigma_{z,x}^{2}$ =$\hat{\sigma}_{z,x}^{2}=\operatorname{Var}(s_{y_{x}}^{(z)}(x))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}||x||^{2}s^{2}\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A17)
$\sigma_{x,z}^{2}$ =$\hat{\sigma}_{x,z}^{2}=\operatorname{Var}(s_{y_{z}}^{(x)}(z))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)||z||^{2}s^{2}\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}||z||^{2}s^{2}\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A18)
$\sigma_{z,z}^{2}$ $=\hat{\sigma}_{z,z}^{2}=\operatorname{Var}(s_{y_{z}}^{(z)}(z))=\frac{1}{S}\left(1-\frac{1}{S}\right)||z||^{2}s^{2}.$ (A19)
Therefore, the differences $\sigma_{x,x}-\sigma_{z,x}$, $\sigma_{x.z}-\sigma_{z,z}$, $\hat{\sigma}_{x,x}-\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{x.z}-\hat{\sigma}_{z,z}$ are negligible for large enough $S$. Hence as long as the simplified model approximates classification tasks to which Lemmas 1 and 2 are applied, these assumptions are reasonably justified.
B.2 Relaxing the assumption of Lemma 1
In Lemma 1 we assume that an attacker has access to the true underlying distribution. However, this is not necessarily the case in real-world settings. Noting that the Equation 12 mainly relies on the true location parameters $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale parameter $\sigma(x)$, we may relax this assumption.
Notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(x)$. From Equation A36 in the proof of Lemma 1 we have
$\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}\end{cases}$ (A20)
$\log\beta=\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}},\end{cases}$ (A21)
where we abuse notations by denoting $\mu_{\mathrm{in}}$ to refer to $\mu_{\mathrm{in}}(x)$ and similarly for other parameters. Since we need to choose a rejection region of the likelihood ratio test such that $\beta\geq 1$, we have
$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A22)
Therefore, the sufficient condition about attacker’s knowledge on the underlying distribution for Lemma 1 to hold is
$\begin{cases}\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A23)
We summarise this discussion in the following:
Lemma 7 (Lemma 1 with a relaxed assumption).
Suppose that the true distribution of $t_{x}$ is of location-scale family with locations $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale $\sigma(x)$, and that LiRA models $t_{x}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x))$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}(x))$. Assume that an attacker estimates parameters $\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x),\hat{\mu}_{\mathrm{out}}(x)$ and $\hat{\sigma}_{\mathrm{out}}(x)$ with an approximated underlying distribution such that
$\begin{cases}\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}(x)+\hat{\mu}_{\mathrm{out}}(x)}{2}+\mu_{\mathrm{out}}(x)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ \sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}(x)+\hat{\mu}_{\mathrm{out}}(x)}{2}+\mu_{\mathrm{out}}(x)\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x).\end{cases}$ (A24)
Then the LiRA vulnerability of a fixed target example is
$\textsc{tpr}_{\mathrm{LiRA}}(x)=\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x),\end{cases}$ (A25)
where $F_{t}$ is the cdf of $t$ with the standard location and unit scale.
Appendix C Missing proofs of Section 3
C.1 Proof of Lemma 1

### v4

@@ 4.2 Experimental results
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1. We provide tabular data and experiments using ResNet-50 in the Appendix (Figures A.1, A3 and A4).
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 2 with tabular data and experiments using ResNet-50 in the Appendix (Figures A.2, A5 and A6). However, the trend w.r.t. $C$ is not as clear as with $S$.
RMIA In Figure 3 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr). More results for RMIA are in Figures A.5, A.6 and A.7 in the Appendix.
@@ 4.3 Model to predict dataset vulnerability
The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3. We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold and Perktold, 2010). The general form of the model can be found in Equation 13, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
@@
We utilise MIA results of ViT-B (Head) (see Table A3) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix. Figure 4 shows the parameters of the prediction model fitted to the training data. For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts.
Figure 4: Coefficient values for different fpr when fitting a regression model based on Equation 13 fitted on data from ViT-B (Head) with LiRA (Table A3). The error bars display the 95% confidence intervals based on Student’s t-distribution. Theoretical values in the simplified model is shown by pink dotted lines ($\beta_{S}=0.5$ and $\beta_{C}=0$).
@@
  • R-50 (FiLM): Figure 5(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section C.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A7 in the Appendix.
@@
(a) The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100). The linear model is robust to changing the feature extractor from ViT-B to R-50.
(b) Regression model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted. 1 left: fine-tuned with FiLM (see Table A7) 2 right: trained from scratch. Data is from Carlini et al. (2022) .
4.4 Individual MIA vulnerability
In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately. In Figure 6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon. The plot shows the maximal vulnerability of all samples and different quantiles over six seeds. The solid line is the median and the errorbars display the min and max over seeds. The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
When fitting the model in Equation 13 we observe that the coefficients $\beta_{S}$ that model the relationship between vulnerability and examples per class for the quantiles are $-0.5603$, $-0.5688$ and $-0.4796$ which is close to the theoretical value of $-0.5$ derived in the theoretical analysis in Section 3.4. However, the maximum vulnerability decreases with a lower slope of $-0.2695$ which is considerably smaller. With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is $-0.3478$ suggesting that higher $S$ is required for the most vulnerable points.
Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon. We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower.
Table 1: The minimum $S$ with $C$=$2$ predicted by the models in Sections 4.3 and 4.4 to empirically match the DP bounds ($\delta$=$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 40054005\,4005 400 69 0006900069\,00069 000 320 000320000320\,000320 000 $5.5\times 10^{9}$ $0.50$ 1 10011001\,1001 100 16 0001600016\,00016 000 88 0008800088\,00088 000 $2.6\times 10^{8}$ $0.75$ $360$ 5 90059005\,9005 900 38 0003800038\,00038 000 $3.5\times 10^{7}$ $1.00$ $160$ 2 70027002\,7002 700 19 0001900019\,00019 000 $7.0\times 10^{6}$
4.5 Comparison between empirical models and universal DP bounds
While the practical evaluation through MIAs is statistical and does not provide universal formal guarantees like DP, the power-law can aid understanding about how the practical vulnerability to MIA behaves in a more realistic threat model when the examples per class increase.
@@ 5 Discussion
Trying to bridge empirical MIA vulnerability and formal DP guarantees is not an easy task because of different threat models and different nature of bounds (statistical vs. universal). While we were able to show both theoretically (Section 3) and empirically (Section 4) that having more examples per class provides protection against MIA in fine-tuned neural networks, the numbers required for significant protection (Section 4.5) limit the practical utility of this observation.
Using a level of formal DP bounds that provide meaningful protection as a yardstick, at least tens of thousands of examples per class are needed for every class even at $\textsc{fpr}=0.001$. The $\epsilon$ values used here are formal upper bounds and must not be compared to actual DP deep learning with comparable privacy budgets, as the latter would be far less vulnerable. At this number of samples per class and a good pre-trained model, the impact of DP training is often negligible. This stresses the importance of formal guarantees like DP for privacy protection.
Broader Impact
Our work systematically studies factors influencing privacy risk of trained ML models. This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
@@
We mostly consider LiRA in our paper which is optimal for our simiplified model of membership inference (Section 3.4) but for the transfer-learning experiments (Section 4) there might be stronger attacks in the future. Furthermore, our results assume well-behaved underlying distributions. Formal bounds on MIA vulnerability would require something like DP. In addition, both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable.
We will publish the code in an open repository after acceptance of the paper.
@@ Acknowledgments
Supplementary Material
Appendix A Details of Section 3
A.1 Formulating LiRA
Let $\mathcal{M}$ be our target model and $\ell(\mathcal{M}(\bm{x}),y)$ be the loss of the model on a target example $(\bm{x},y)$. The goal of MIA is to determine whether $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$. This can be formulated as a hypothesis test:
$H_{0}:(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$ (A1)
$H_{1}:(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}.$ (A2)
Following (Carlini et al., 2022), we formulate the Likelihood Ratio Attack (LiRA). LiRA exploits the difference of losses on the target model under $H_{0}$ and $H_{1}$. To model the IN/OUT loss distributions with few shadow models, LiRA employs a parametric modelling. Particularly, LiRA models $t_{\bm{x}}$ by a normal distribution. That is, the hypothesis test formulated above can be rewritten as
$H_{0}^{\prime}:t_{\bm{x}}\sim\mathcal{N}(\hat{\mu}_{\mathrm{out}},\hat{\sigma}_{\mathrm{out}}^{2})$ (A3)
$H_{1}^{\prime}:t_{\bm{x}}\sim\mathcal{N}(\hat{\mu}_{\mathrm{in}},\hat{\sigma}_{\mathrm{in}}^{2}).$ (A4)
@@ A.1 Formulating LiRA
$\mathrm{LR}(\bm{x})=\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{in}},\hat{\sigma}_{\mathrm{in}}^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{out}},\hat{\sigma}_{\mathrm{out}}^{2})}.$ (A5)
LiRA rejects $H_{0}^{\prime}$ if and only if
$\mathrm{LR}(\bm{x})\geq\tau,$ (A6)
concluding that $H_{1}^{\prime}$ is true, i.e., identifying the membership of $(\bm{x},y)$. Thus, the true positive rate of this hypothesis test given as
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\mathrm{LR}(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (A7)
@@
A.2 On the assumption of shared scale
@@
In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$. Using the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),...,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Then the IN/OUT LiRA scores are given as
$s_{y}^{\mathrm{(in)}}$ $=\langle\bm{x},\frac{1}{S}\left(\sum_{i=1}^{S-1}\bm{x}_{y,i}+\bm{x}\right)\rangle=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle+\langle\bm{x},\frac{1}{S}(\bm{x}-\bm{x}_{y,S})\rangle$ (A8)
$s_{y}^{\mathrm{(out)}}$ $=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle.$ (A9)
@@
Since for the simplified model scores follow Gaussian distributions, $\sigma_{\mathrm{in}}=\hat{\sigma}_{\mathrm{in}}$ and $\sigma_{\mathrm{out}}=\hat{\sigma}_{\mathrm{out}}$. It follows that
@@
$\sigma_{\mathrm{in}}^{2}$ $=\hat{\sigma}_{\mathrm{in}}^{2}=\operatorname{Var}(s_{y}^{\mathrm{(in)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)-\frac{1}{S^{2}}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\left(1-\frac{1}{S}\right)\bm{x}^{T}\Sigma\bm{x}$ (A10)
$\sigma_{\mathrm{out}}^{2}$ $=\hat{\sigma}_{\mathrm{out}}^{2}=\operatorname{Var}(s_{y}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\bm{x}^{T}\Sigma\bm{x}.$ (A11)
Thus, the differences $\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}-\hat{\sigma}_{\mathrm{out}}$ are negligible for large $S$.
A.3 Proof of Lemma 1